OpenStudy (anonymous):

Draw a line labeled 2 y equals minus x minus 1 by joining the ordered pairs negative 3,1 and 3, negative 2. Draw a line labeled 2 y equals x plus 5 by joining the ordered pairs negative 3, 1 and 3, 4 he x-coordinate of the solution to the system of equations is _____.

OpenStudy (anonymous):

@DivineSolar

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

the link doesnt do anything, all they want you to do from the looks of the description, is plot a couple points and connect them with a line

OpenStudy (danjs):

2y = -x-1 points (-3 , 1) and (3,-2)

OpenStudy (anonymous):

actulley this is the quetion for that graph The x-coordinate of the solution to the system of equations is _____.

OpenStudy (danjs):



OpenStudy (anonymous):

so 2y=-x -1 is the answer

OpenStudy (danjs):

the x coordinate of the solution to the system, is the x coordinate of the point (x,y) where the two lines intersect

OpenStudy (danjs):

no, that is just the first line, you have to graph the second one too

OpenStudy (anonymous):

oh sorry im bad at math

OpenStudy (danjs):

The 2 lines are 2y = -x - 1 2y = x+5 since both are already solved for 2y = ..... you can set both the right sides equal to find the x value of the solution -x - 1 = x + 5 solve for x

OpenStudy (anonymous):

sorry im not sure how too

OpenStudy (danjs):

or graph both of them like they want you to, and look at the point where they intersect (x,y) the x value is the answer...

OpenStudy (anonymous):

i really dont know how to do this sorry

OpenStudy (danjs):